# Ask, Reason, Assist: Decentralized Robot Collaboration via Language and Logic

### v1

@@ Abstract
Increased robot deployment, such as in warehousing, has revealed a need for seamless collaboration among heterogeneous robot teams to resolve unforeseen conflicts. To address this challenge, we propose a novel decentralized framework that enables robots to request and provide help. The process begins when a robot detects a conflict and uses a Large Language Model (LLM) to decide whether external assistance is required. If so, it crafts and broadcasts a natural language (NL) help request. Potential helper robots reason over the request and respond with offers of assistance, including information about the effect on their ongoing tasks. Helper reasoning is implemented via an LLM grounded in Signal Temporal Logic (STL) using a Backus–Naur Form (BNF) grammar, ensuring syntactically valid NL-to-STL translations, which are then solved as a Mixed Integer Linear Program (MILP). Finally, the requester robot selects a helper by reasoning over the expected increase in system-level total task completion time. We evaluated our framework through experiments comparing different helper-selection strategies and found that considering multiple offers allows the requester to minimize added makespan. Our approach significantly outperforms heuristics such as selecting the nearest available candidate helper robot, and achieves performance comparable to a centralized “Oracle” baseline but without heavy information demands.
@@ I Introduction
Modern warehouses and manufacturing sites are complex ecosystems where heterogeneous multi-robot teams must coordinate to meet demanding throughput targets. However, the dynamic nature that makes these mixed fleets effective also renders them brittle to unforeseen physical and semantic conflicts. Given the scale of modern-day robot deployments, a decentralized yet efficient conflict-resolution process is essential to unlock the full potential of these teams.
Recent advances in foundation models, particularly multi-modal Large Language Models (LLMs), provides a promising tool in addressing the first step of this process - identifying and describing the conflict [1]. Paired with an embodied agent, by processing sensor inputs and an understanding of robot capabilities, these models can generate a concise Help Request in Natural Language [2]. Although methods have been presented to use multi modal LLMs for end-to-end conflict resolution in multi-robot settings [3], the plans generated by foundation models do not provide the safety guarantees required in environments where humans and robots coexist. Furthermore, industrial robots frequently operate under strict time constraints, such as in order fulfillment scenarios where shipments must be staged at outbound docks by designated pickup times. Without guarantees on the safety and feasibility of conflict recovery plans, LLMs alone are insufficient for ensuring reliable multi-agent conflict resolution in time-sensitive and safety-critical environments.
One path forward to solving this challenge is to apply formal methods, such as the verification of Signal Temporal Logic (STL) and Linear Temporal Logic (LTL) specifications. STL and LTL tools can guarantee that a robot’s plan satisfies constraints and have been successfully applied to multi-robot coordination and reconfiguration tasks [4][5]. Recent work has also explored their integration with LLMs to enforce safety constraints during task planning [6]. However, it remains open how to integrate these formal methods approaches with foundation model reasoning in a way that enables seamless, safe multi-agent coordination.
@@ I-A Related Work
I-A1 Formal Methods for Task and Motion Planning
Formal methods, such as Temporal Logic (TL), provide a precise language for specifying robot objectives with temporal and spatial grounding. In particular, Signal Temporal Logic (STL) is well-suited to robotics as it reasons over real-valued trajectories with explicit time bounds [7]. In task and motion planning (TAMP), these requirements are often formulated as a mixed-integer program (MIP) which unifies discrete task choices and continuous motion in one certifiable solver [8]. Our framework builds directly on this paradigm but differs on where and how formal reasoning is invoked. Once an LLM translates NL to STL on demand, each potential helper independently solves a compact MILP that (i) certifies feasibility and (ii) estimates the marginal cost of adding the help task given its existing task schedule that remains to be private. This parallelized approach maintains the rigorous guarantee provided by formal methods while avoiding the need for a centralized planner and full disclosure of potentially proprietary information.
@@ I-A2 Natural Language to Temporal Logic
While temporal logic is a powerful formal language for expressing constraints on task and motion plans, it requires significant effort and expertise to formulate. As such, a number of works have sought to automatically translate natural language (NL) instructions and specifications into temporal logic (TL) formulas, specifically using LLMs [9, 10, 11, 12, 13]. However, they lack guarantees on the accuracy and syntactic validity of generated temporal logic formulas. Recently, conformal prediction has been used to quantify the uncertainty of LLM outputs, and applied towards an accuracy guarantee for NL to TL [12]. Specifically, an LLM is prompted to incrementally translate an NL specification, with the LLM queried multiple times for each translation step. Then, the frequency of each unique response at each step is used as a measure of uncertainty, and triggers asking for help from a human when high uncertainty is detected. Syntactic validity in this work is addressed by repeatedly sampling from the LLM with the same prompt and filtering out responses which contain invalid operators or atomic propositions. We note that there is no guarantee that any of the sampled responses from the LLM will be syntactically valid, and that these filtering rules do not fully capture the syntax rules of temporal logic. We seek to address this problem in this work.
@@ I-A3 Meta-heuristics in Vehicle Routing Problems
As the centralized baseline for comparison, assigning a new help request within a fleet can be formulated as a dynamic Vehicle Routing Problem (VRP), a long studied subject within the operations research field [14]. Since exact optimization is intractable at realistic scales, constructive insertion heuristics (e.g sequential/greedy insertion [15]) provide a feasible solution promptly and meta-heuristics like Iterated Local Search (ILS) improve the optimality of the initial solution by introducing local moves with perturbations to escape local minima [16]. We implement this centralized hybrid VRP algorithm (insertion + ILS) as an “Oracle” baseline that can reshuffle tasks across all robots to minimize the fleet’s overall makespan. In contrast to this centralized VRP approach, our framework only requires compact exchanges in NL $\rightarrow$ TL $\rightarrow$ MILP bids, minimizing the information demand without sacrificing scalability.
Contributions
We present a framework that enables heterogeneous warehouse robots, each subject to local task specifications, to request and offer assistance in natural language, while ensuring the safety and feasibility of the resulting recovery plan through formal verification. In particular, we make the following contributions:
@@
    We propose a method to transform natural language specifications into temporal logic specifications with a guarantee on validity (Secs. III-A and III-B).
@@
    We rigorously evaluate our decentralized method against both simple heuristics and a near-optimal centralized “Oracle” baseline demonstrating a compelling trade-off between performance and information demand. (Sec. IV-B).
@@ II Preliminaries and Problem Statement
Consider robots operating in parallel on a variety of tasks; in this work we use warehousing as our canonical example. We investigate how a system can achieve decentralized conflict resolution while approaching the global efficiency of a centralized solution through peer-to-peer interactions that minimizes the information demand. If a robot detects a conflict that it cannot resolve on its own, how can it request and receive help from a fellow robot while ensuring safety and minimizing impact on the overall system? We now formalize these notions.
@@
We consider a multi-robot system (MRS) with $n$ robots, and $\mathcal{R}=\{0,1,\cdots,n-1\}$ denoting the set of all robots. The robots operate in a discrete time grid world, so robot $i$ is at $x_{i}(t)\in W\subset\mathbb{Z}^{2}$ at time $t=0,1,\cdots,H$, where $H\in\mathbb{N}$ is a finite time horizon. Robot $i$’s trajectory is a sequence of its position from time $t\in[0,H]$, denoted by
$X_{i}=\{x_{i}(t)\}_{t=0}^{H}\subseteq W.$ (1)
We index a trajectory as $X_{i}(j)=x_{i}(j)$ at a specific timestep and $X_{i}(j:k)=\{x_{i}(t)\}_{t=j}^{k}$ for a range. Given trajectory $X_{i}$ of duration $H$, its Manhattan distance is
$D(X_{i})=\sum_{t=0}^{H-1}\|x_{i}(t+1)-x_{i}(t)\|_{1}.$ (2)
Each cell is either free space or a static obstacle, a common setup in modeling multi-robot systems [17]; we denote all free cells as $F\subseteq W$. Multiple robots can occupy a free cell concurrently ($x_{i}(t)=x_{j}(t)\in F$ is allowed); we assume local coordination and collision avoidance are handled by lower-level planning and control. Finally, each robot has an NL representation of its own capabilities $c_{i}\in\mathbb{L}$ (e.g., “can lift pallets” or “max speed $1~$\mathrm{m}\mathrm{/}\mathrm{s}$\$”)
@@
For each robot $i$, its assigned tasks $\Psi_{i}$ are represented as an STL specification $\varphi_{i}$ and path $X_{i}$ for robot $i$. All robots must also obey a global specification [^1] See example in Sec. -A1 for actuation limits and obstacles $\varphi^{\mathrm{\textnormal{g}}}$. We assume each robot’s tasks are feasible while, we certify the syntactic validity of the STL via BNF constrained generation (Sec. III-B).
@@
Our goal is for robots to minimize time to complete their own tasks plus help tasks, which is the makespan of each robot and the overall system. Consider robot $i$, with initial position $x_{i}(0)\in F$ and corresponding STL specification $\varphi_{i}$. If the robot’s trajectory $X_{i}$ satisfies $\varphi_{i}$, we write $X_{i}\models\varphi_{i}$. Note, this expression is a shorthand for $(X_{i},0)\models\varphi_{i}$, meaning we always evaluate the satisfaction of the trajectory starting from $t=0$. We minimize makespan by finding $X_{i}$ to take the least time to satisfy $\varphi_{i}$. To aid in defining the makespan we first define the time to first satisfaction $T(X_{i},\varphi)_{i}$ as
@@
That is, the makespan finds a trajectory that minimizes the time to first satisfaction of $\varphi_{i}$ while obeying the robot’s initial condition. We implement 4 as an MILP solved with Gurobi [18]. Further details are in Sec. IV.
@@
We assume robots can send and receive natural language messages $m\in\mathbb{L}$ instantaneously and error-free. We denote $m_{i\to j}$ as a message from robot $i$ to robot $j$, and a broadcast message as $m_{i\to\mathcal{R}}$.
@@
Suppose robot $i$ detects a conflict $(m^{\mathrm{\textnormal{c}}},x^{\mathrm{\textnormal{c}}},\tau^{\mathrm{\textnormal{c}}})$. Without using centralized task assignment, we seek to identify another robot $j\in\mathcal{R}_{-i}=\mathcal{R}\setminus\{i\}$ is the set of candidate helper robots, and plan its motion such that it resolves the conflict while minimizing the increase in the total makespan across all robots. That is, create $\varphi_{j}$ such that $x_{j}(t)=x^{\mathrm{\textnormal{c}}}$ for all $t^{\prime}\in[t,t+\tau^{\mathrm{\textnormal{c}}}]$ while minimizing the sum of individual makespans $\sum_{i\in\mathcal{R}}M(x_{i}(0),\varphi_{i})$.
@@ III Proposed Method
In this section, we introduce the main components of our framework, as depicted in Fig. LABEL:fig:overall_framework. Concretely, Sec. III-A describes how a requester robot that requires help generates a help request and broadcasts it to the multi-robot system. Each robot in the system then evaluates its capabilities and availability, and any helper robot with the potential to assist formulates a help offer in natural language (NL). Then, Sec. III-B details how each helper translates its help request from NL to STL, while Sec. III-C describes how each helper computes an optimal path plan that minimizes total time impact to the system. Each potential helper then replies to the requester in natural language with its help offer and the associated, predicted time impact to the system. Finally, the requester chooses the helper with the lowest impact and affirms its selection via a help confirmation message.
@@ III-A Generating Help Requests, Offers, and Confirmations
We implement a multi-robot communication protocol for collaborative conflict resolution using LLMs. This protocol involves three capabilities: send, receive and broadcast. We define three types of messages: help requests, help offers, and help confirmations [^2] An example help request is in Sec. -B1 and help offer in Sec. -B2..
Help requests are broadcast messages sent by a robot facing a conflict after determining that it requires help, by reasoning over the conflict $(m^{\mathrm{\textnormal{c}}},x^{\mathrm{\textnormal{c}}},\tau^{\mathrm{\textnormal{c}}})$ and its own capabilities $c_{i}$, conditioned on a prompt $p^{\mathrm{\textnormal{c}}}\in\mathbb{L}$:
@@
Help offers are sent by potential helpers in response to a help request. To generate help offers, each robot $j$ reasons over the help request, its own location, and its capabilities:
@@
We again use constrained generation to ensure help offers include information about capabilities. In addition, the help offer includes the duration it will take robot $j$ to provide help $\tau^{\mathrm{\textnormal{h}}}_{j}$ (i.e., how long robot $i$ needs to wait), and the additional time to help relative to completing its original tasks $\tau^{\mathrm{\textnormal{new}}}_{j}$ (i.e., how much robot $j$ is impacted by having to complete the help request). We compute these durations below in Sec. III-C (see 12).
Finally, the requester confirms the lowest-cost help offer by finding

### v3

# Ask, Reason, Assist: Robot Collaboration via Natural Language and Temporal Logic

@@ Abstract
Increased robot deployment, such as in warehousing, has revealed a need for collaboration among heterogeneous robot teams to resolve unforeseen conflicts. To this end, we propose a peer-to-peer coordination protocol that enables robots to request and provide help without a central task allocator. The process begins when a robot detects a conflict and uses a Large Language Model (LLM) to decide whether external assistance is required. If so, it crafts and broadcasts a natural language (NL) help request. Potential helper robots reason over the request and respond with offers of assistance, including information about the effect on their ongoing tasks. Helper reasoning is implemented via an LLM grounded in Signal Temporal Logic (STL) using a Backus–Naur Form (BNF) grammar, ensuring syntactically valid NL-to-STL translations, which are then solved as a Mixed Integer Linear Program (MILP). Finally, the requester robot selects a helper by reasoning over the expected increase in system-level total task completion time. We evaluated our framework through experiments comparing different helper-selection strategies and found that considering multiple offers allows the requester to minimize added makespan. Our approach significantly outperforms heuristics such as selecting the nearest available candidate helper robot, and achieves performance comparable to a centralized “Oracle” baseline but without heavy information demands.
@@ I Introduction
Modern warehouses deploy heterogeneous robot fleets, combining mobile robots, forklifts, and manipulators, to meet demanding throughput targets. The diversity of capabilities that makes these teams effective also introduces coordination challenges: robots encounter physical conflicts (e.g., blocked paths) and semantic conflicts (e.g., incompatible task types) that a single robot cannot resolve on its own. At scale, centralized conflict resolution becomes impractical and may require disclosing proprietary schedules. Realizing the full potential of these teams requires an efficient peer-to-peer approach where robots request and provide help without revealing private task information.
Recent advances in foundation models, particularly multimodal Large Language Models (LLMs), provides a promising tool in addressing the first step of the conflict resolution process: identifying and describing the conflict [20]. When paired with an embodied agent, these models can generate a concise help request in natural language by processing sensor inputs while leveraging an understanding of robot capabilities [4]. Although methods have been presented to use multimodal LLMs for end-to-end conflict resolution in multi-robot settings [16], the plans generated by foundation models do not provide the safety guarantees required in environments where humans and robots coexist,nor the spatio-temporal context needed to compute feasible paths under timing constraints. Furthermore, industrial robots frequently operate under strict time constraints, such as in order fulfillment scenarios where shipments must be staged at outbound docks by designated pickup times. Without guarantees on the safety and feasibility of conflict recovery plans, LLMs alone are insufficient for ensuring reliable multi-agent conflict resolution in time-sensitive and safety-critical environments.
One path forward to solving this challenge is to apply formal methods, such as the verification of Signal Temporal Logic (STL) and Linear Temporal Logic (LTL) specifications. STL and LTL tools can guarantee that a robot’s plan satisfies constraints and have been successfully applied to multi-robot coordination and reconfiguration tasks [13][23]. Recent work has also explored their integration with LLMs to enforce safety constraints during task planning [28]. However, it remains open how to integrate these formal methods with foundation model reasoning for multi-agent coordination. A key insight is that transmitting formal specifications directly between robots is impractical in heterogeneous fleets: each robot has its own atomic propositions grounded in its capabilities; a STL specification composed in one robot’s vocabulary may not be interpreted by another robot without a separate integration layer. Natural language bridges this gap, allowing the requester to describe what is needed while each helper determines how to encode it in its own specification.
@@ I-A Related Work
I-A1 Formal Methods for Robot Coordination
Formal methods, such as Temporal Logic (TL), provide a precise language for specifying robot objectives with temporal and spatial grounding. In particular, Signal Temporal Logic (STL) is well-suited to robotics as it reasons over real-valued trajectories with explicit time bounds [1]. Prior work formulates these specifications as mixed-integer programs (MIPs) for certifiable planning [17]. We adopt this approach as a local solver at each robot, and contribute a coordination protocol for robots to exchange help requests and leverage their private solvers without central oversight. That is, our work differs on where and how formal reasoning is invoked. Once an LLM translates NL to STL on demand, each potential helper independently solves a compact Mixed Integer Linear Program (MILP) that (i) certifies feasibility and (ii) estimates thee marginal cost of adding the help task given its existing, private task schedule. This parallelized approach maintains the rigorous guarantee provided by formal methods while avoiding the need for a centralized planner and full disclosure of potentially proprietary information.
@@ I-A2 Natural Language to Temporal Logic
While temporal logic is a powerful formal language for expressing constraints on task and motion plans, it requires significant effort and expertise to formulate. As such, a number of works have sought to automatically translate natural language (NL) instructions and specifications into temporal logic (TL) formulas, specifically using LLMs [5, 12, 19, 27, 10]. However, they lack guarantees on the accuracy and syntactic validity of generated temporal logic formulas. Recently, conformal prediction has been used to quantify the uncertainty of LLM outputs, and applied towards an accuracy guarantee for NL to TL [27]. Specifically, an LLM is prompted to incrementally translate an NL specification, with the LLM queried multiple times for each translation step. Then, the frequency of each unique response at each step is used as a measure of uncertainty, and triggers asking for help from a human when high uncertainty is detected. Syntactic validity in this work is addressed by repeatedly sampling from the LLM with the same prompt and filtering out responses which contain invalid operators or atomic propositions. We note that there is no guarantee that any of the sampled responses from the LLM will be syntactically valid, and that these filtering rules do not fully capture the syntax rules of temporal logic. We seek to address the problem of generating correct temporal logic syntax in this work.
@@ I-A3 Meta-heuristics in Vehicle Routing Problems
As the centralized baseline for comparison, assigning a new help request within a fleet can be formulated as a dynamic Vehicle Routing Problem (VRP), a long-studied subject within the operations research field [7]. Since exact optimization is intractable at realistic scales, constructive insertion heuristics (e.g sequential/greedy insertion [21]) provide a feasible solution promptly and meta-heuristics like Iterated Local Search (ILS) improve the optimality of the initial solution by introducing local moves with perturbations to escape local minima [18]. We implement this centralized hybrid VRP algorithm (insertion + ILS) as an “Oracle” baseline that can reshuffle tasks across all robots to minimize the fleet’s overall makespan. In contrast to this centralized VRP approach, our framework requires only brief NL messages and scalar cost bids, where robots never disclose their full schedules, therefore minimizing information exchange without sacrificing scalability.
I-B Contributions
We present a framework for heterogeneous warehouse robots, each subject to local task specifications, to request and offer assistance in NL, while ensuring the safety and feasibility of the resulting recovery plan through formal verification. In particular, we make the following contributions:
@@ I-B Contributions
    We propose a method to transform NL specifications into TL specifications with a guarantee on syntactic validity (Secs. III-A and III-B).
@@
    We rigorously evaluate our coordination protocol against both simple heuristics and a centralized “Oracle” baseline, demonstrating competitive performance with minimal information exchange. (Sec. IV-B).
@@ II Preliminaries and Problem Statement
Consider robots operating in parallel on a variety of tasks; in this work we use warehousing as our canonical example. We investigate how a system can achieve decentralized conflict resolution while approaching the global efficiency of a centralized solution through peer-to-peer interactions that minimizes the information demand. Specifically, we ask: If a robot detects a conflict that it cannot resolve on its own, how can it request and receive help from a fellow robot while ensuring safety and minimizing impact, measured as the added time cost, on the overall system?
@@
[…]
Each cell is either free space or a static obstacle, a common setup in modeling multi-robot systems [22]; we denote all free cells as $F\subseteq W$. Multiple robots can occupy a free cell concurrently ($x_{i}(t)=x_{j}(t)\in F$ is allowed); we assume local coordination and collision avoidance are handled by lower-level planning and control for which many solutions exist (e.g., [24, 25, 11]). Finally, each robot has an NL representation of its own capabilities $c_{i}\in\mathbb{L}$ (e.g., “can lift pallets” or “max speed $1~$\mathrm{m}\mathrm{/}\mathrm{s}$\$”).
@@
For each robot $i$, its assigned tasks $\Psi_{i}$ are represented as an STL specification $\varphi_{i}$ and path $X_{i}$ for robot $i$. All robots must also obey a global specification $\varphi^{\mathrm{\textnormal{g}}}$; see App.-A1 for an example including actuation limits and obstacles. We assume each robot’s tasks are feasible, while we certify syntactic validity via constrained generation (Sec. III-B).
@@
Our goal is for robots to minimize time to complete their own tasks plus help tasks, which is the makespan of each robot and the overall system. Consider robot $i$, with initial position $x_{i}(0)\in F$ and corresponding STL specification $\varphi_{i}$. If the robot’s trajectory $X_{i}$ satisfies $\varphi_{i}$, we write $X_{i}\models\varphi_{i}$. Note, this expression is a shorthand for $(X_{i},0)\models\varphi_{i}$, meaning we always evaluate the satisfaction of the trajectory starting from $t=0$. We minimize makespan by finding $X_{i}$ to take the least time to satisfy $\varphi_{i}$. To aid in defining the makespan, define the time to first satisfaction
@@
That is, the makespan finds a trajectory that minimizes the time to first satisfaction of $\varphi_{i}$ while obeying the robot’s initial condition. We implement 2 as an MILP solved with Gurobi [14]. Further details are in Sec. IV.
@@
We assume robots can send and receive NL messages $m\in\mathbb{L}$ instantaneously and error-free. We use NL rather than transmitting STL specification $\varphi_{i}$ directly, as valid STL generation requires receiver-side context such as its atomic propositions. We denote $m_{i\to j}$ as a message from robot $i$ to robot $j$, and a broadcast message as $m_{i\to\mathcal{R}}$.
@@
We now formalize the research problem using the above definitions: Suppose robot $i$ detects a conflict $(m^{\mathrm{\textnormal{c}}},x^{\mathrm{\textnormal{c}}},\tau^{\mathrm{\textnormal{c}}})$. Without using centralized task assignment, we seek to identify another robot $j\in\mathcal{R}_{-i}=\mathcal{R}\setminus\{i\}$ among the set of candidate helper robots, and plan its motion such that it resolves the conflict while minimizing the increase in the total makespan across all robots. That is, create $\varphi_{j}$ such that $x_{j}(t)=x^{\mathrm{\textnormal{c}}}$ for all $t^{\prime}\in[t,t+\tau^{\mathrm{\textnormal{c}}}]$ while minimizing the sum of individual makespans $\sum_{i\in\mathcal{R}}M(x_{i}(0),\varphi_{i})$.
@@ III Proposed Method
In this section, we introduce the main components of our framework, as depicted in Fig. LABEL:fig:overall_framework. Concretely, Sec. III-A describes how a requester robot that requires help generates a help request and broadcasts it to the multi-robot system. Each robot in the system then evaluates its capabilities and availability, and any helper robot with the potential to assist formulates a help offer in natural language (NL). Then, Sec. III-B details how each helper translates its help request from NL to STL, while Sec. III-C describes how each helper computes an optimal path that minimizes total time impact to the system. Each potential helper then replies to the requester in natural language with its help offer and the associated, predicted time impact to the system. Finally, the requester chooses the helper with the lowest impact and affirms its selection via a help confirmation message.
@@ III-A Generating Help Requests, Offers, and Confirmations
We implement a multi-robot communication protocol for collaborative conflict resolution using LLMs. This protocol involves three capabilities: send, receive and broadcast. We define three types of messages: help requests, help offers, and help confirmations. See examples in App.-B1 and -B2.
Help requests are broadcast messages sent by a robot $i$ facing a conflict after determining that it requires help, by reasoning over the conflict $(m^{\mathrm{\textnormal{c}}},x^{\mathrm{\textnormal{c}}},\tau^{\mathrm{\textnormal{c}}})$ and its own capabilities $c_{i}$, conditioned on a prompt $p^{\mathrm{\textnormal{c}}}\in\mathbb{L}$:
@@
Help offers are sent by potential helpers in response to a help request. Each helper $j$ generates a help offer by independently grounding the NL request into its own STL specification by reasoning over the request, its location, capabilities, and its task schedule, which remains private:
@@
We again use constrained generation to ensure help offers include information about capabilities. In addition, the help offer includes the duration it will take robot $j$ to provide help $\tau^{\mathrm{\textnormal{h}}}_{j}$ (i.e., how long robot $i$ needs to wait), and the additional time to help relative to completing its original tasks $\tau^{\mathrm{\textnormal{new}}}_{j}$ (i.e., how much robot $j$ is impacted by having to complete the help request). We compute these durations below in Sec. III-C (see 10). Finally, the requester confirms the lowest-cost help offer: